# DINO-Foresight: Looking into the Future with DINO

### v1

@@ 5 Conclusion
We introduced DINO-Foresight, a future prediction framework that forecasts VFM features instead of raw RGB frames. Our encoder-decoder transformer architecture efficiently propagates multi-layer, high-resolution VFM features, enabling strong performance across a range of scene understanding tasks. By leveraging these features, the masked transformer predicts future state representations for multiple tasks, including semantic segmentation, instance segmentation, depth estimation, and surface normal prediction. Extensive evaluations show our method’s excellent performance, with intermediate transformer features further enhancing results, underscoring the potential of our approach for self-supervised visual representation learning.
Future work could explore scaling our method with larger, more diverse datasets and incorporating stronger VFMs like DINOv2-Huge or multiple VFMs with varied scene understanding capabilities. Additionally, exploring action conditioning to enhance controllability in autonomous driving is another promising direction for future research.
@@ Acknowledgements
This work has been partially supported by project MIS 5154714 of the National Recovery and Resilience Plan Greece 2.0 funded by the European Union under the NextGenerationEU Program.
Hardware resources were granted with the support of GRNET. Also, this work was performed using HPC resources from GENCI-IDRIS (Grants 2023-A0141014182, 2023-AD011012884R2, and 2024-AD011012884R3).
@@ Appendix A Additional Results
A.1 Impact of PCA
In Table 6, we evaluate the impact of applying PCA to reduce the dimensionality of the multi-layer VFM features. We see that using PCA improves the semantic segmentation forecasting results for short-term prediction and especially of moving objects. However, on mid-term prediction, the differences are eliminated. In depth forecasting, PCA enhances performance across all metrics for both short-term and mid-term predictions.
Method Segmentation Depth
@@ A.1 Impact of PCA
Without PCA 71.3 70.4 59.9 57.6 87.9 .122 84.8 .147
With PCA 71.8 71.7 59.8 57.6 88.6 .114 85.4 .136
Table 6: Impact of PCA. Results on semantic segmentation and depth forecasting.
Method Segmentation Depth Surface Normals
@@
Oracle 77.0 77.4 77.0 77.4 89.1 .108 89.1 .108 3.24 95.3 3.24 95.3
Copy Last 54.7 52.0 40.4 32.3 84.1 .154 77.8 .212 4.41 89.2 5.39 84.0
VISTA [27] 45.7 43.8 40.4 35.6 84.9 .143 80.2 .183 4.87 85.9 5.63 83.2
DINO-Foresight 71.8 71.7 59.8 57.6 88.6 .114 85.4 .136 3.39 94.4 4.00 91.3
Table 7: Comparison across tasks. We use DINOv2 encoder and show performance on segmentation (ALL, MO), depth estimation ($\delta_{1}$ accuracy, AbsRel error), and surface normal prediction (m, percentage within 11.25°). For the VISTA baseline, we used DepthAnythingV2 [94] and Lotus [35] to generate depth and surface normal predictions respectively.
A.2 More Visualizations
@@
In Figs. 5 and 6, we present additional qualitative results illustrating the prediction of semantic segmentation, depth maps, and surface normals. Specifically, we compare our method, DINO-Foresight, against the Oracle, which involves using future RGB frames as inputs for different prediction heads, as well as Vista [27]. For Vista, the outputs are derived from the Segmenter [76], DepthAnythingV2 [94], and Lotus [35], applied to RGB frames predicted by Vista. In these two scenes, Vista performs well in capturing static elements when the vehicle is moving forward, as illustrated in Fig. 5. However, its quality diminishes when dealing with moving objects and smaller structures, as shown in Fig. 6. In both cases, DINO-Foresight excels at maintaining the integrity of motion dynamics and accurately capturing intricate details.
@@
In Figs. 7 and 6, we offer additional qualitative results derived from utilizing DINO-Foresight for the prediction of semantic segmentation and depth maps and surface normals over extended time intervals. These outcomes are achieved through the use of autoregressive rollouts. Beginning with a series of four context frames ($X_{t-9}$ to $X_{t}$), the model is capable of predicting up to 48 subsequent frames, equivalent to 2.88 seconds, with predictions occurring at an interval of every third frame. Our model consistently delivers accurate predictions over the entire forecasted duration, effectively capturing motion dynamics and maintaining consistency across different modalities. This performance underscores its robustness and versatility, which are related to its capability of predicting the features of a foundation model. As a final remark, it is important to note that the noisy segmentation predictions observed at the bottom of the images, present in both the predicted and Oracle results, are attributed to unannotated regions in the Cityscapes dataset that are disregarded during DPT training. This artifact impacts only the segmentation outcomes of DPT head and does not affect the predicted future features, as evidenced by the clear and accurate depth and surface normal predictions.
@@
Figure 5: Visualization of future predictions for semantic segmentation, depth, and surface normals. The illustrated scene is Frankfurt (01 (017082-017111)). Both DINO-Foresight and Vista capture the static elements effectively when the vehicle is moving forward, highlighting their capabilities in such scenarios. Noisy segmentation predictions at the bottom of the image (in both predicted and Oracle results) are due to unannotated regions in Cityscapes that are ignored during DPT training. This artifact affects only segmentation, not the predicted future features, as evident in the clear depth and surface normal predictions.
@@
Figure 6: Visualization of future predictions for semantic segmentation, depth, and surface normals. The illustrated scene is Munster (15). Vista struggles with moving objects and smaller structures, whereas DINO-Foresight maintains more detailed predictions. Noisy segmentation predictions at the bottom of the image (in both predicted and Oracle results) are due to unannotated regions in Cityscapes that are ignored during DPT training. This artifact affects only segmentation, not the predicted future features, as evident in the clear depth and surface normal predictions.
@@
Figure 7: Long-term semantic segmentation, depth and surface normal predictions. The illustrated scene is Frankfurt (01 (011791-011820)). DINO-Foresight consistently preserves motion dynamics and intricate details in complex scenes over extended time horizons. Noisy segmentation predictions at the bottom of the image (in both predicted and Oracle results) are due to unannotated regions in Cityscapes that are ignored during DPT training. This artifact affects only segmentation, not the predicted future features, as evident in the clear depth and surface normal predictions.
@@
Figure 8: Long-term semantic segmentation, depth and surface normal predictions. The illustrated scene is Frankfurt (01 (006570-006599)). DINO-Foresight excels in predicting the motion of the nearby car but faces challenges with distant, low-motion objects, highlighting areas for future improvement. Noisy segmentation predictions at the bottom of the image (in both predicted and Oracle results) are due to unannotated regions in Cityscapes that are ignored during DPT training. This artifact affects only segmentation, not the predicted future features, as evident in the clear depth and surface normal predictions.
Appendix B Limitations and Future Work
In this work, we introduced DINO-Foresight, a simple yet effective method for semantic future prediction based on forecasting VFM features. While the results are promising, several exciting research opportunities remain.
First, our approach uses a simple masked transformer architecture with a SmoothL1 loss, which is deterministic and does not account for the inherent uncertainty in future predictions. Incorporating stochastic elements could better capture this ambiguity, leading to more robust predictions. Although we explored strategies to reduce training compute for high-resolution feature prediction, inference-time compute demands remain unchanged. Future work could address this by adopting hierarchical transformer architectures [72], which would not only improve efficiency but also enable the model to handle even higher feature resolutions.
To reduce the dimensionality of multi-layer VFM features, we relied on PCA, a linear projection method. While effective, exploring advanced dimensionality reduction techniques, such as VAEs, could further enhance feature representation and boost performance.
Another promising direction lies in leveraging the simple and scalable design of DINO-Foresight to investigate its behavior with larger datasets and bigger models. Our strong results on Cityscapes suggest that scaling both the data and model size, including the VFM encoder and prediction transformer, could unlock further improvements in future forecasting.
Finally, extending DINO-Foresight to incorporate action conditioning and action prediction presents an exciting opportunity to transform it into a world model capable of reasoning and control.
Appendix C Implementation Details of Prediction Heads
We provide implementation details for the heads trained on different downstream tasks. The DPT head is used for semantic segmentation, depth estimation, and surface normal estimation. We adopt the DPT [70] implementation from Depth Anything [93, 94], setting the feature dimensionality to 256 and configuring dptoutchannels = [128, 256, 512, 512]. For all tasks, models are trained for 100 epochs with a batch size of 128 ($16\times 8$ GPUs). The learning rate is set to 0.0016, using the AdamW optimizer with linear warmup for the first 10 epochs, and weight decay is 0.0001. For semantic segmentation, we use a polynomial scheduler and cross-entropy loss with 19 classes. For depth estimation, we use a cosine annealing scheduler and cross-entropy loss, with 256 classes. For surface normal estimation, we employ a polynomial scheduler and a loss function combining cosine similarity and $L_{2}$ loss with weighted averaging, using 3 classes.
For the Mask2Former head used in instance segmentation, we implement our approach using the official Mask2Former [17] and Detectron2 [90] codebases. The main difference compared to the official Mask2Former configuration for Cityscapes instance segmentation is the input feature maps. In our approach, the four multi-scale feature maps expected by Mask2Former are derived from the PCA features. These PCA features are first projected to 128, 256, 512, and 1024 dimensions and then resized so their spatial resolutions are ×$\times 4$, ×$\times 2$, ×$\times 1$, and ×$\times 0.5$ relative to the original resolution of the DINOv2 ViT-B outputs. We train using the AdamW optimizer, with a batch size of 64 ($8\times 8$ GPUs), learning rate of 0.00032, weight decay of 0.05, and 67,500 iterations, with a polynomial scheduler.

### v2

@@ 5 Conclusion
In this work, we introduced DINO-Foresight, a self-supervised framework for semantic future prediction that shifts the paradigm from forecasting low-level latent representations to predicting high-dimensional, semantically rich VFM features. This shift offers several key advantages: it enhances scene understanding by leveraging structured semantic information, avoids the complexity of full-frame synthesis, and enables scalable, modular integration with downstream tasks.
To realize this approach, we designed a masked feature transformer that efficiently propagates high-resolution, multi-layer VFM features over time. Our experiments show that forecasting such features is not only feasible but also highly effective—demonstrating strong performance across diverse future-frame understanding tasks including semantic segmentation, instance segmentation, depth estimation, and surface normal prediction. Unlike prior methods that rely on multiple task-specific models, our single model handles all tasks seamlessly, validating the scalability and versatility of our framework.
Overall, this work lays the foundation for a new class of unified and modular future prediction systems, grounded in semantic reasoning rather than low-level reconstruction.
@@ Acknowledgements
This work has been partially supported by project MIS 5154714 of the National Recovery and Resilience Plan Greece 2.0 funded by the European Union under the NextGenerationEU Program. Hardware resources were granted with the support of GRNET. Also, this work was performed using HPC resources from GENCI-IDRIS (Grants 2023-A0141014182, 2023-AD011012884R2, and 2024-AD011012884R3).
Appendix
@@ Appendix A Additional Results
A.1 Impact of Dimensionality Reduction
In our work, we examine a PCA-based dimensionality reduction method and find that compressing features in this way does not compromise performance on semantic segmentation and depth prediction downstream tasks (Table 5). In fact, reducing the dimensionality simplifies the modeling process and leads to improved performance. Specifically, for semantic segmentation forecasting, PCA enhances short-term predictions—particularly for moving objects—while its effect on mid-term predictions is negligible. Similarly, for depth forecasting, dimensionality reduction consistently boosts performance across all metrics for both short- and mid-term predictions.
Table 5: Impact of Dimensionality Reduction. Reduction is performed using PCA. Results on semantic segmentation and depth forecasting.
Dim. Reduction Segmentation Depth
@@ A.1 Impact of Dimensionality Reduction
71.3 70.4 59.9 57.6 87.9 .122 84.8 .147
71.8 71.7 59.8 57.6 88.6 .114 85.4 .136
A.2 Emerging Visual Representations in the Future-Frame Masked Feature Transformer
Self-supervised representation learning has achieved remarkable progress, with numerous studies focusing on extracting robust visual features from unlabeled images and videos (Bardes et al., 2024; Tong et al., 2022; Girdhar et al., 2023; Ryali et al., 2023; Wang et al., 2023; Chen et al., 2020; Caron et al., 2021; He et al., 2022; Grill et al., 2020; Wei et al., 2022; Kakogeorgiou et al., 2024, 2022; Gidaris et al., 2024; Venkataramanan et al., 2025; Sirko-Galouchenko et al., 2025). Inspired by these advancements, we investigate the potential of our future-frame masked feature transformer as a self-supervised method for enhancing VFM visual features. Specifically, we train DPT heads for semantic segmentation and depth prediction, using not only the features predicted by the masked feature transformer but also additional features extracted from intermediate transformer layers. We examine features from the 6th, 9th, 10th, 11th, and 12th (last) layers of the transformer to assess whether these intermediate representations can further improve the strong VFM features predicted by our masked transformer.
Results, shown in Figure 4, indicate that incorporating intermediate transformer features from the masked transformer enhances segmentation performance, with one exception (6th-layer features for short-term segmentation). Notably, the best segmentation results are achieved using features from the 9th layer. Although the improvements are modest, this aligns with expectations given the strength of the predicted VFM features alone. Similar results are observed for future depth prediction, where intermediate features also led to performance gains, with the best depth results achieved using 6th-layer features (see Figure 4). While exploring self-supervised learning was not the primary aim of our work, we find these results intriguing, as they suggest that future prediction methods hold promise as self-supervised visual representation learners. We hope this work can spark further research into this direction.
A.3 Additional Comparisons with VISTA
In Table 6, we present a comprehensive evaluation comparing our method against VISTA (Gao et al., 2024) for semantic segmentation, instance segmentation, depth and surface normal estimation on the Cityscapes dataset. We extend this evaluation to the nuScenes (Caesar et al., 2020) dataset in Table 7 for depth and surface normals. Notably, DINO-Foresight demonstrates strong zero-shot generalization capabilities. When trained only on Cityscapes and directly evaluated on nuScenes without fine-tuning, performance degradation is minimal, while still significantly outperforming all baselines. The results show that DINO-Foresight consistently surpasses the performance of VISTA on both datasets.
Table 6: Comparison across tasks on Cityscapes. We use DINOv2 encoder and show performance on segmentation (ALL, MO), instance segmentation (AP50, AP), depth estimation ($\delta_{1}$ accuracy, AbsRel error), and surface normal prediction (m, percentage within 11.25°). VISTAft is the VISTA model fine-tuned on Cityscapes.
Method Semantic Segm. Instance Segm Depth Surface Normals
Short Mid Short Mid Short Mid Short Mid
[…]
Oracle 77.0 77.4 77.0 77.4 66.2 40.4 66.2 40.4 89.1 .108 89.1 .108 3.24 95.3 3.24 95.3
Copy Last 54.7 52.0 40.4 32.3 24.7 10.4 9.5 2.8 84.1 .154 77.8 .212 4.41 89.2 5.39 84.0
VISTAft 64.9 62.1 53.9 51.0 33.1 17.7 19.8 9.0 86.4 .124 82.8 .153 3.75 93.0 4.30 90.0
DINO-Foresight 71.8 71.7 59.8 57.6 50.5 26.6 27.3 12.6 88.6 .114 85.4 .136 3.39 94.4 4.00 91.3
Table 7: Comparison across tasks at nuScenes. We use DINOv2 encoder and show performance on depth estimation ($\delta_{1}$ accuracy, AbsRel error) and surface normal prediction (m, percentage within 11.25°). VISTAft is the VISTA model fine-tuned on nuScenes. The last row evaluates zero-shot generalization: DINO-Foresight trained on Cityscapes and directly evaluated on nuScenes.
Method Depth Surface Normals
Mid Long Mid Long
[…]
Oracle 82.6 .206 82.6 .206 3.09 97.1 3.09 97.1
Copy Last 73.4 .353 68.4 .468 4.66 88.6 5.31 85.3
VISTAft 74.6 .337 70.8 .421 4.46 90.8 4.96 88.2
DINO-Foresight 80.7 .218 76.3 .299 3.59 93.9 4.22 90.6
DINO-Foresight (zero-shot) 78.4 .269 72.1 .377 4.03 92.3 4.77 88.4
A.4 Impact of Multi-Layer Features
Scene understanding models benefit from utilizing features from multiple layers of a frozen image encoder. To fully exploit the pretrained DINO features, we integrate representations from several layers into the DPT head. Table 8 presents an ablation comparing multi-layer DINO features to using only the final layer (layer 12) for semantic segmentation on Cityscapes. The results demonstrate that aggregating features from layers 3, 6, 9, and 12 enhances performance, with ALL (all semantic classes) and MO (movable objects classes) scores rising from 72.1/73.4 to 77.0/77.4.
Table 8: DINO+DPT Features Ablation. Ablation of multi-layer DINO features as input to the DPT head versus using only the last layer features, evaluated on semantic segmentation on Cityscapes.
Layers Segmentation
ALL MO
12 72.1 73.4
3,6,9,12 77.0 77.4
A.5 Masking strategy and loss function ablations
We evaluate the impact of masking strategies and loss functions on forecasting performance. As shown in Table 9, full masking (masking all future features) consistently outperforms random masking across all metrics, for semantic segmentation and depth forecasting. This demonstrates that masking all future features forces the model to learn more robust temporal dynamics. Regarding loss functions (Table 10), we find that our framework is robust to loss function choice, with L1, MSE, SmoothL1, and SmoothL1+Cosine achieving comparable performance.
Table 9: Impact of Masking Strategies on Cityscapes. We compare random masking versus full masking for semantic segmentation, depth estimation, and surface normal prediction. Full masking demonstrates consistently superior performance across all metrics.
Masking Strategy Semantic Segmentation Depth Surface Normals
@@ A.5 Masking strategy and loss function ablations
Random Masking 70.5 70.2 58.0 55.7 88.2 .121 84.7 .148 3.48 94.0 4.11 90.8
Full Masking 71.8 71.7 59.8 57.6 88.6 .114 85.4 .136 3.39 94.4 4.00 91.3
Table 10: Loss Function Comparison on Cityscapes. We evaluate different loss functions (L1, MSE, SmoothL1, SmoothL1+Cosine) for semantic segmentation, depth estimation, and surface normal prediction. Results demonstrate that our framework is robust to loss function choice, with all variants achieving comparable performance.
Loss Function Semantic Segmentation Depth Surface Normals
Short Mid Short Mid Short Mid
[…]
L1 71.7 71.7 59.7 57.6 88.6 .118 85.6 .138 3.41 94.3 4.00 91.3
MSE 71.7 71.9 60.0 57.8 88.6 .117 85.4 .138 3.40 94.4 3.99 91.4
SmoothL1+Cos 71.7 71.4 59.8 57.5 88.7 .116 85.5 .137 3.40 94.4 3.98 91.4
SmoothL1 71.8 71.7 59.8 57.6 88.6 .114 85.4 .136 3.39 94.4 4.00 91.3
A.6 Model Size and Data Scale Scalability
We investigate how performance scales with model size and training data diversity. As shown in Table 11, we evaluate three model variants—Small (115M), Base (258M), and Large (460M) parameters—by modifying the hidden dimension and number of attention heads while keeping dataset size and training duration fixed. Results demonstrate consistent performance improvements with increased model capacity, particularly for mid-term predictions, indicating that larger models better capture complex temporal dynamics in VFM features. Regarding data scale (Table 12), we combine Cityscapes and nuScenes datasets with equal-probability sampling during training and evaluate on Cityscapes. The model trained on combined datasets achieves consistent improvements across all tasks compared to training on Cityscapes alone, with gains particularly pronounced for mid-term semantic segmentation. These findings demonstrate that our framework effectively scales with both increased model capacity and diverse training data, motivating further exploration with larger models and datasets to fully realize the potential of forecasting VFM features for multi-task scene understanding.
Table 11: Model Size Scalability on Cityscapes. We evaluate three model sizes—Small (115M), Base (258M), and Large (460M) parameters—across semantic segmentation, depth estimation, and surface normal prediction. Results demonstrate consistent performance improvements with increased model capacity.
Model Variant Hidden Dim Att Heads Semantic Segm. Depth Surface Normals
Short Mid Short Mid Short Mid
[…]
Small (115M) 768 6 71.1 70.8 59.3 57.3 87.7 .125 84.8 .142 3.58 93.7 4.13 90.8
Base (258M) 1152 8 71.8 71.7 59.8 57.6 88.6 .114 85.4 .136 3.39 94.4 4.00 91.3
Large (460M) 1536 12 71.9 71.7 60.2 58.3 88.6 .116 85.5 .137 3.40 94.4 4.00 91.5
Table 12: Data Scale Scalability. We compare training on Cityscapes alone versus training on combined Cityscapes+nuScenes data. Results demonstrate that increasing training data diversity leads to improved performance across all tasks and temporal horizons.
Training Data Semantic Segmentation Depth Surface Normals
Short Mid Short Mid Short Mid
[…]
Cityscapes 71.8 71.7 59.8 57.6 88.6 .114 85.4 .136 3.39 94.4 4.00 91.3
Cityscapes+nuScenes 72.3 72.2 61.0 59.4 88.6 .117 85.7 .136 3.41 94.4 3.95 91.6
A.7 More Visualizations
@@
In Figure 5 and 6, we present additional qualitative results illustrating the prediction of semantic segmentation, depth maps, and surface normals. Specifically, we compare our method, DINO-Foresight, against the Oracle, which involves using future RGB frames as inputs for different prediction heads, as well as VISTA Gao et al. (2024).
As illustrated in Figure 5, DINO-Foresight maintains superior integrity of motion dynamics and geometric consistency across frames, resulting in more accurate predictions
@@
In Figure 7 and 8, we offer additional qualitative results derived from utilizing DINO-Foresight for the prediction of semantic segmentation and depth maps and surface normals over extended time intervals. These outcomes are achieved through the use of autoregressive rollouts. Beginning with a series of four context frames ($X_{t-9}$ to $X_{t}$), the model is capable of predicting up to 48 subsequent frames, equivalent to 2.88 seconds, with predictions occurring at an interval of every third frame. Our model consistently delivers accurate predictions over the entire forecasted duration, effectively capturing motion dynamics and maintaining consistency across different modalities. This performance underscores its robustness and versatility, which are related to its capability of predicting the features of a foundation model. As a final remark, it is important to note that the noisy segmentation predictions observed at the bottom of the images, present in both the predicted and Oracle results, are attributed to unannotated regions in the Cityscapes dataset that are disregarded during DPT training. This artifact impacts only the segmentation outcomes of DPT head and does not affect the predicted future features, as evidenced by the clear and accurate depth and surface normal predictions.
Appendix B Limitations and Future Work
In this work, we introduced DINO-Foresight, a simple yet effective method for semantic future prediction based on forecasting VFM features. Our approach delivers strong results while opening several exciting directions for future research.
First, our method uses a straightforward masked transformer with SmoothL1 loss. While forecasting VFM features avoids the challenges of modeling complex pixel distributions, our current implementation is deterministic. However, our framework can easily be extended to capture uncertainty—for example, by adding a diffusion loss (as in MAR (Li et al., 2024)) or modeling tokens with a Gaussian mixture model (as in GIVT (Tschannen et al., 2024)). These extensions would better handle future ambiguity while maintaining the simplicity of our approach.
Although we explored strategies to reduce training compute for high-resolution feature prediction, inference-time compute demands remain unchanged. Future work could address this by adopting hierarchical transformer architectures (Ryali et al., 2023), which would not only improve efficiency but also enable the model to handle even higher feature resolutions.
Another promising direction is scaling DINO-Foresight to larger datasets and models. Our experiments on model size (115M to 460M parameters) and data diversity (Cityscapes+nuScenes) demonstrate consistent performance improvements, particularly for mid-term predictions, suggesting that further scaling to even larger models and more diverse datasets could yield substantial gains in forecasting performance.. Furthermore, the flexibility of our framework allows seamless integration of newer VFM encoders, such as RADIOv2.5 (Heinrich et al., 2025), which combines multiple VFMs into a single, more powerful model, enhancing its multi-task future scene understanding capabilities.
Appendix C Broader Impact
Our work enables efficient and scalable semantic future prediction by forecasting semantically rich VFM features. This allows flexible integration with different scene understanding tasks without retraining, making it useful for applications like autonomous driving and robotics. While we do not foresee risks in our approach, we must remain mindful that the pretrained Vision Foundation Models we build upon may carry biases, which could potentially influence our semantic future predictions.
Appendix D Implementation Details
We provide implementation details for the heads trained on different downstream tasks. The DPT head is used for semantic segmentation, depth estimation, and surface normal estimation. We adopt the DPT Ranftl et al. (2021) implementation from Depth Anything Yang et al. (2024a, b), setting the feature dimensionality to 256 and configuring dptoutchannels = [128, 256, 512, 512]. For all tasks, models are trained for 100 epochs with a batch size of 128 ($16\times 8$ GPUs). The learning rate is set to 0.0016, using the AdamW optimizer with linear warmup for the first 10 epochs, and weight decay is 0.0001. For semantic segmentation, we use a polynomial scheduler and cross-entropy loss with 19 classes. For depth estimation, we use a cosine annealing scheduler and cross-entropy loss, with 256 classes. For surface normal estimation, we employ a polynomial scheduler and a loss function combining cosine similarity and $L_{2}$ loss with weighted averaging, using 3 classes.
For the Mask2Former head used in instance segmentation, we implement our approach using the official Mask2Former Cheng et al. (2022) and Detectron2 Wu et al. (2019) codebases. The main difference compared to the official Mask2Former configuration for Cityscapes instance segmentation is the input feature maps. In our approach, the four multi-scale feature maps expected by Mask2Former are derived from the PCA features. These PCA features are first projected to 128, 256, 512, and 1024 dimensions and then resized so their spatial resolutions are $\times 4$, $\times 2$, $\times 1$, and $\times 0.5$ relative to the original resolution of the DINOv2 ViT-B outputs. We train using the AdamW optimizer, with a batch size of 64 ($8\times 8$ GPUs), learning rate of 0.00032, weight decay of 0.05, and 67,500 iterations, with a polynomial scheduler.
Regarding Vista, we fine-tuned the model with 8 frames in total (3 as context and 5 future frames) for cityscapes, while for Nuscenes we used 9 frames in total (3 as context and 6 future frames) to support long-term forecasting.
Appendix E Definitions of Evaluation Metrics
[…]
Appendix F Details of Evaluation Scenarios
Regarding Cityscapes, the target frame for both short and mid term prediction is 20. We subsample sequences by a factor of 3 before inputting to the model. For short-term prediction, the model uses frames 8, 11, 14, and 17 as context to predict frame 20 (with context length $N_{c}=4$ and $N_{p}=1$). For mid-term prediction, the model uses frames 2, 5, 8, and 11 as context and predicts frame 20 auto-regressively through frames 14 and 17. We calculate segmentation metrics on the 20th frame using Cityscapes ground truth. For depth and surface normals, we rely on pseudo-annotations from DepthAnythingV2 Yang et al. (2024b) and Lotus He et al. (2025), respectively, due to the lack of true annotations in Cityscapes. Regarding nuScenes, the target frame for both mid and long term prediction is 29. Again, we subsample sequences by a factor of 3 before input to the model. For mid-term prediction, the model uses frames 11, 14, 17, and 20 as context and predicts frame 29 auto-regressively through frames 23 and 26. For long-term prediction, the model uses frames 2, 5, 8, and 11 as context and predicts frame 29 auto-regressively through frames 14,17,20,23 and 26. Again for depth and surface normals, we rely on pseudo-annotations from DepthAnythingV2 Yang et al. (2024b) and Lotus He et al. (2025), respectively, due to the lack of true annotations in nuScenes.
@@ Appendix F Details of Evaluation Scenarios
Figure 5: Visualization of future predictions for semantic segmentation, depth, and surface normals. The illustrated scene is Frankfurt (01 (017082-017111)).
@@
Figure 6: Visualization of future predictions for semantic segmentation, depth, and surface normals. The illustrated scene is Munster (15).
@@
Figure 7: Long-term semantic segmentation, depth and surface normal predictions. The illustrated scene is Frankfurt (01 (011791-011820)). DINO-Foresight consistently preserves motion dynamics and intricate details in complex scenes over extended time horizons.
@@
Figure 8: Long-term semantic segmentation, depth and surface normal predictions. The illustrated scene is Frankfurt (01 (006570-006599)). DINO-Foresight excels in predicting the motion of the nearby car but faces challenges with distant, low-motion objects, highlighting areas for future improvement.